# Sponsored Search Auction Design Beyond Single Utility Maximization

### v1

@@ 4.2 Large Market Auction
corollary 1 gives a bounded approximation for the scenario where $\max_{i\in[n]}v_{i}\cdot\alpha_{1}\geq\frac{\mathrm{OPT}}{\rho}$. This subsection addresses the converse scenario – the large market case, where $\max_{i\in[n]}v_{i}\cdot\alpha_{1}<\frac{\mathrm{OPT}}{\rho}$. In a large market, no bidder or slot dominates. In addition, we also assume, w.l.o.g., that $\rho$ is sufficiently large, e.g. $\rho\geq 4$. Consequently, we can leverage the random sampling technique to develop a truthful mechanism with a bounded approximation ratio.
@@
Let $w_{1}\geq w_{2}\geq...\geq w_{\ell}$ be positive real numbers, such that the sum $w=w_{1}+...+w_{\ell}$ satisfies $w_{1}<w/4$. We independently partition each number with equal probability into two subsets $A$ and $B$. With probability at least $1/2$, there exists a matching $\mathcal{M}$ between $A$ and $B$ such that $\sum_{(w_{i},w_{j})\in\mathcal{M}}\min\{w_{i},w_{j}\}>\frac{w}{4}$.
@@
$\sum_{(w_{j},w_{j})\in\mathcal{M}}\min\{w_{i},w_{j}\}\geq\frac{w}{4}$
@@
Consider a bad event in ${\cal B}$ and the corresponding partition $(A,B)$. Let $k$ denote the size of the smaller set among the two subsets. We match the $t$-th largest number in $A$ to the $t$-th largest number in $B$ for each index $t\in[k]$, obtaining a matching $\mathcal{M}$ with size $k$. Furthermore, let $\bar{A}\subseteq A$ and $\bar{B}\subseteq B$ represent the smaller elements in each pair of the matching (breaking ties arbitrarily). Due to the definition of a bad event, we have
$\sum_{(w_{i},w_{j})\in\mathcal{M}}\min\{w_{i},w_{j}\}=\sum_{w_{i}\in\bar{A}}w_{i}+\sum_{w_{j}\in\bar{B}}w_{j}\leq\frac{w}{4}.$
Define a set $C$ as the numbers that are not in $\bar{A}\cup B$, i.e., $C:=(A\setminus\bar{A})\cup(B\setminus\bar{B})$. Sort the elements in $C$ from largest to smallest, and let $C_{1}$ and $C_{2}$ respectively denote the sets formed by the numbers positioned at odd and even indices. A new event $(A^{\prime},B^{\prime})$ is constructed by letting $A^{\prime}=C_{1}\cup\bar{A}$ and $B^{\prime}=C_{2}\cup\bar{B}$. Clearly, for each bad event, such a new event $(A^{\prime},B^{\prime})$ is unique.
@@
$\sum_{w_{i}\in C_{1}}w_{j}+\sum_{w_{j}\in C_{2}}w_{j}=w-(\sum_{w_{i}\in\bar{A}}w_{i}+\sum_{w_{j}\in\bar{B}}w_{j})\geq\frac{3v}{4}$
and the large market assumption $w_{1}<w/4$, we have
$\sum_{(w_{i},w_{j})\in\mathcal{M}^{\prime}}\min\{w_{i},w_{j}\}>\frac{1}{2}\cdot(\frac{3w}{4}-\frac{v}{4})=\frac{w}{4},$
@@
Intuitively, lemma 3 asserts that with high probability, there exists a matching between $k$ slots and $R$ such that if for each slot $j\in[k]$, we enforce the matched bidder in $R$ purchases it at a unit price of $z_{j}$ once her bid is no less than $z_{j}$, the total obtained payment is at least $\frac{1}{4}$ of the optimal social welfare. However, the mechanism cannot enforce such actions as it would lead to untruthfulness. Therefore, we employ a technique of discounted sales to ensure both truthfulness and a constant approximation ratio simultaneously. We show the following theorem.
@@
Under the large market assumption that $\max_{i\in[n]}v_{i}\cdot\alpha_{1}<\frac{\mathrm{OPT}}{4}$,  algorithm 3 is truthful and obtains an approximation ratio at most $64$.
@@
Now we analyze the approximation ratio. Let $\Pi^{*}=\{\pi^{*}_{i}\}_{i\in[n]}$ be an optimal assignment. Viewing $b_{i}\alpha_{\pi^{*}_{i}}$ as $w_{i}$ in lemma 3, we have with probability at least $1/2$, there exists a matching $\mathcal{M}$ between $S^{(1)}$ and $S^{(2)}$ such that $\sum_{(i,h)\in\mathcal{M}}\min\{b_{i}\alpha_{\pi^{*}_{i}},b_{h}\alpha_{\pi^{*}_{h}}\}>\mathrm{OPT}/4.$ Refer to such a partition $\{S^{(1)},S^{(2)}\}$ as a good partition. The following proof is conditioned on the occurrence of a good partition and analyzes the conditional expected performance of our algorithm. For brevity, we omit the notation for conditional expectations and use $\mathbb{E}[\cdot]$ to denote the conditional expectation.
@@
$\mathbb{E}[{\cal P}({\cal A})]\geq\frac{1}{2}\cdot\sum_{(i,h)\in\mathcal{M}}\min\{b_{i}\alpha_{\pi^{*}_{i}},b_{h}\alpha_{\pi^{*}_{h}}\}>\frac{\mathrm{OPT}}{8}.$ (6)
@@
By combining equation 6, equation 7, and the probability of a good partition occurring, we can demonstrate that the expected approximation ratio of algorithm 3 is at most $64$ under the large market assumption.
@@
3:   With probability of $1/17$ begin
@@
6:   With probability of $16/17$ begin
@@
$\mathbb{E}[\mathrm{ALG}]=\frac{1}{17}\cdot\mathrm{ALG}_{1}+\frac{16}{17}\cdot\mathrm{ALG}_{2}.$
When the instance satisfies the large market assumption, due to theorem 4.3,
$\mathbb{E}[\mathrm{ALG}_{2}]\geq\frac{1}{64}\cdot\mathrm{OPT};$
@@
$\mathrm{ALG}_{1}\geq\frac{1}{4(1+\epsilon)}\cdot\mathrm{OPT}.$
Combining the above two cases proves that  algorithm 4 is a $68(1+\epsilon)$-approximation mechanism.

### v2

@@ 4.2 Large Market Auction
corollary 1 gives a bounded approximation for the scenario where $\max_{i\in[n]}v_{i}\cdot\alpha_{1}\geq\frac{\mathrm{OPT}}{\rho}$. This subsection addresses the converse scenario – the large market case, where $\max_{i\in[n]}v_{i}\cdot\alpha_{1}<\frac{\mathrm{OPT}}{\rho}$. In a large market, no bidder or slot dominates. Consequently, we can leverage the random sampling technique to develop a truthful mechanism with a bounded approximation ratio.
@@
Let $w_{1}\geq w_{2}\geq...\geq w_{\ell}$ be positive real numbers, such that the sum $w=w_{1}+...+w_{\ell}$ satisfies $w_{1}<w/\rho$. We independently partition each number with equal probability into two subsets $A$ and $B$. With probability at least $1/2$, there exists a matching $\mathcal{M}$ between $A$ and $B$ such that
$\sum_{(w_{i},w_{j})\in\mathcal{M}}\min\{w_{i},w_{j}\}>\frac{w}{3}\cdot\left(1-\frac{1}{\rho}\right).$
@@
$\sum_{(w_{j},w_{j})\in\mathcal{M}}\min\{w_{i},w_{j}\}\geq\frac{w}{3}\cdot\left(1-\frac{1}{\rho}\right)$
@@
Consider a bad event in ${\cal B}$ and the corresponding partition $(A,B)$. Let $k$ denote the size of the smaller set among the two subsets. We match the $t$-th largest number $a_{t}$ in $A$ to the $t$-th largest number $b_{t}$ in $B$ for each index $t\in[k]$, obtaining a matching $\mathcal{M}$ with size $k$. Furthermore, we define $\bar{A}:=\{a_{t}\in A|a_{t}\leq b_{t}\}$, which represents the elements in $A$ that are the minimum in their respective matching pairs. Similarly, we define $\bar{B}:=\{b_{t}\in B|b_{t}<a_{t}\}$. Due to the definition of a bad event, we have
$\sum_{(w_{i},w_{j})\in\mathcal{M}}\min\{w_{i},w_{j}\}=\sum_{w_{i}\in\bar{A}}w_{i}+\sum_{w_{j}\in\bar{B}}w_{j}\leq\frac{w}{3}\cdot\left(1-\frac{1}{\rho}\right).$
Define a set $C$ as the numbers that are not in $\bar{A}\cup\bar{B}$, i.e., $C:=(A\setminus\bar{A})\cup(B\setminus\bar{B})$. Sort the elements in $C$ from largest to smallest, and let $C_{1}$ and $C_{2}$ respectively denote the sets formed by the numbers positioned at odd and even indices. A new event $(A^{\prime},B^{\prime})$ is constructed by letting $A^{\prime}=C_{1}\cup\bar{A}$ and $B^{\prime}=C_{2}\cup\bar{B}$. Clearly, for each bad event, such a new event $(A^{\prime},B^{\prime})$ is unique.
@@
$\sum_{w_{i}\in C_{1}}w_{j}+\sum_{w_{j}\in C_{2}}w_{j}=w-(\sum_{w_{i}\in\bar{A}}w_{i}+\sum_{w_{j}\in\bar{B}}w_{j})\geq\left(\frac{2}{3}+\frac{1}{3\rho}\right)\cdot w$
and the large market assumption $w_{1}<w/\rho$, we have
$\sum_{(w_{i},w_{j})\in\mathcal{M}^{\prime}}\min\{w_{i},w_{j}\}>\frac{w}{2}\cdot\left(\frac{2}{3}+\frac{1}{3\rho}-\frac{1}{\rho}\right)=\frac{w}{3}\cdot\left(1-\frac{1}{\rho}\right),$
@@
Intuitively, lemma 3 asserts that with high probability, there exists a matching between $k$ slots and $R$ such that if for each slot $j\in[k]$, we enforce the matched bidder in $R$ purchases it at a unit price of $z_{j}$ once her bid is no less than $z_{j}$, the total obtained payment is at least $\frac{1}{3}\cdot\left(1-\frac{1}{\rho}\right)$ of the optimal social welfare. However, the mechanism cannot enforce such actions as it would lead to untruthfulness. Therefore, we employ a technique of discounted sales to ensure both truthfulness and a constant approximation ratio simultaneously. We show the following theorem.
@@
Under the large market assumption that $\max_{i\in[n]}v_{i}\cdot\alpha_{1}<\frac{\mathrm{OPT}}{\rho}$,  algorithm 3 is truthful and obtains an approximation ratio at most $48/(1-1/\rho)$.
@@
Now we analyze the approximation ratio. Let $\Pi^{*}=\{\pi^{*}_{i}\}_{i\in[n]}$ be an optimal assignment. Viewing $b_{i}\alpha_{\pi^{*}_{i}}$ as $w_{i}$ in lemma 3, we have with probability at least $1/2$, there exists a matching $\mathcal{M}$ between $S^{(1)}$ and $S^{(2)}$ such that
$\sum_{(i,h)\in\mathcal{M}}\min\{b_{i}\alpha_{\pi^{*}_{i}},b_{h}\alpha_{\pi^{*}_{h}}\}>\frac{\mathrm{OPT}}{3}\cdot\left(1-\frac{1}{\rho}\right)~{}.$
Refer to such a partition $\{S^{(1)},S^{(2)}\}$ as a good partition. The following proof is conditioned on the occurrence of a good partition and analyzes the conditional expected performance of our algorithm. For brevity, we omit the notation for conditional expectations and use $\mathbb{E}[\cdot]$ to denote the conditional expectation.
@@
$\mathbb{E}[{\cal P}({\cal A})]\geq\frac{1}{2}\cdot\sum_{(i,h)\in\mathcal{M}}\min\{b_{i}\alpha_{\pi^{*}_{i}},b_{h}\alpha_{\pi^{*}_{h}}\}>\frac{\mathrm{OPT}}{6}\cdot\left(1-\frac{1}{\rho}\right).$ (6)
@@
By combining equation 6, equation 7, and the probability of a good partition occurring, we can demonstrate that the expected approximation ratio of algorithm 3 is at most $48/(1-1/\rho)$ under the large market assumption.
@@
3:   With probability of $\sqrt{3}/(12+\sqrt{3})$ begin
@@
6:   With probability of $12/(12+\sqrt{3})$ begin
@@
$\mathbb{E}[\mathrm{ALG}]=\frac{\sqrt{3}}{12+\sqrt{3}}\cdot\mathrm{ALG}_{1}+\frac{12}{12+\sqrt{3}}\cdot\mathrm{ALG}_{2}.$
When the instance satisfies the large market assumption parameterized by $\rho=4\sqrt{3}+1$, due to theorem 4.3,
$\mathbb{E}[\mathrm{ALG}_{2}]\geq\frac{\sqrt{3}}{12+48\sqrt{3}}\cdot\mathrm{OPT};$
@@
$\mathrm{ALG}_{1}\geq\frac{1}{(4\sqrt{3}+1)(1+\epsilon)}\cdot\mathrm{OPT}.$
Combining the above two cases proves that  algorithm 4 admits an approximation ratio of $(49+8\sqrt{3})(1+\epsilon)\approx 62.856$.